### v1

Abstract
Nonsmooth nonconvex-concave minimax problems have attracted significant attention due to their wide applications in many fields. In this paper, we consider a class of nonsmooth nonconvex-concave minimax problems on Riemannian manifolds. Owing to the nonsmoothness of the objective function, existing minimax manifold optimization methods cannot be directly applied to solve this problem. We propose a manifold proximal gradient descent ascent (MPGDA) algorithm for solving the problem. At each iteration, the proposed algorithm alternately performs one or multiple manifold proximal gradient descent steps and a proximal ascent step. We prove that the MPGDA algorithm can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point of the considered problem within $\mathcal{O}(\varepsilon^{-3})$ iterations. Numerical experiments on fair sparse PCA and sparse spectral clustering are conducted to demonstrate the advantages of the MPGDA algorithm.
Keywords:
nonconvex-concave minimax problem nonsmooth manifold optimization manifold proximal gradient descent ascent iteration complexity
MSC:
90C26 90C30 65K05
1 Introduction
In this paper, we consider a class of nonsmooth nonconvex-concave minimax optimization problems formulated as follows:
$\min_{x\in\mathcal{M}}\max_{y\in S}F(x,y):=f(x,y)+h(x)-g(y),$ (1)
where $\mathcal{M}$ is a Riemannian manifold embedded in a finite-dimensional Euclidean space $\mathbb{E}_{1}$, and $S$ is a convex set in another finite-dimensioanl Euclidean space $\mathbb{E}_{2}$, $f:\mathbb{E}_{1}\times\mathbb{E}_{2}\to\mathbb{R}$ is a continuously differentiable function and is concave with respect to y, $h:\mathbb{E}_{1}\to\mathbb{R}$ and $g:\mathbb{E}_{2}\to\mathbb{R}$ are some proper closed convex (possibly nonsmooth) functions. Throughout this paper, we adopt the following assumptions for problem (1) .
{assumption}
$(i)$ The manifold $\mathcal{M}$ is compact.
$(ii)$ The function $f$ has Lipschitz continuous gradient with respect to $x$, i.e., there exists a Lipschitz constant $L_{x}>0$ such that
$\left\|\nabla_{x}f(x_{1},y)-\nabla_{x}f(x_{2},y)\right\|\leq L_{x}\left\|x_{1}-x_{2}\right\|$
for any $x_{1},x_{2}\in\mathcal{M}$ and $y\in S$.
$(iii)$ The function $\nabla_{y}f(x,y)$ is Lipschitz continuous with respect to $(x,y)$, i.e., there exists a Lipschitz constant $L_{y}>0$ such that
$\left\|\nabla_{y}f(x_{1},y_{1})-\nabla_{y}f(x_{2},y_{2})\right\|^{2}\leq L_{y}^{2}(\left\|x_{1}-x_{2}\right\|^{2}+\left\|y_{1}-y_{2}\right\|^{2})$
for any $(x_{1},y_{1})$, $(x_{2},y_{2})\in\mathcal{M}\times S$.
$(iv)$ The function $h$ is $L_{h}$ Lipschitz continuous.
Problem (1) has many important applications in machine learning and signal processing. Below, we present two representative examples and refer the interested readers to chenProximalGradientMethod2020 ; BriefIntroductiontoManifoldOptimization for more examples.
Example 1.
Fair Sparse Principle Component Analysis (FSPCA). The classical PCA Hotelling1933Analysis is one of the most widely used dimensional reduction techniques, which minimizes the total reconstruction error over the whole samples. However, in real-life applications, the data samples come from diverse classes and PCA may exhibit a higher reconstruction error for certain classes than the others. To address this disparity, the fair PCA is proposed samadiPriceFairPCA . Suppose the $m$ data samples belong to $n$ classes, and each class $i$ corresponds to $A_{i}\in\mathbb{R}^{m_{i}\times d}$ with $d$ being the dimension of the samples, $m_{i}$ being the number of samples in the $i$-th class, and $\sum_{i=1}^{n}m_{i}=m$. Then the fair PCA can be formulated as:
$\min_{X\in\mathrm{St}(d,r)}\max_{i=1,\dots,n}-\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X),$ (2)
where $r<d$ is the number of principle components, $\mathrm{Tr}(\cdot)$ represents the trace and $\mathrm{St}(d,r):=\{X\in\mathbb{R}^{d\times r}|X^{\top}X=I_{r}\}$ with $I_{r}$ denoting the $r\times r$ identity matrix. Recently, the fair sparse PCA is further proposed to promote sparsity in the principle components babuFairPrincipalComponent2023 as follows:
$\min_{X\in\mathrm{St}(d,r)}\max_{i=1,\dots,n}-\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X)+\mu\left\|X\right\|_{1},$ (3)
where $\mu>0$ is a weighting parameter and $\left\|X\right\|_{1}:=\sum_{i,j}\left|x_{ij}\right|$ is the $\ell_{1}$-norm of the matrix $X$. In view of (Beck2017First, , Example 4.10), problem (3) can be equivalently reformulated into
$\min_{X\in\mathrm{St}(d,r)}\max_{y\in\Delta_{n}}-\sum_{i=1}^{n}y_{i}\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X)+\mu\left\|X\right\|_{1}.$ (4)
where $\Delta_{n}:=\{y|\sum_{i=1}^{n}y_{i}=1,y_{i}\geq 0,i=1,\dots,n\}$ is the standard simplex in $\mathbb{R}^{n}$. Note that problem (4) is an instance of problem (1) with $\mathcal{M}$ being the Stiefel manifold $\mathrm{St}(d,r)$, $S$ being the standard simplex $\Delta_{n}$, $h(X)=\mu\|X\|_{1}$ and $f(X,y)=-\sum_{i=1}^{n}y_{i}\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X)$.
Example 2.
Sparse Spectral Clustering (SSC). The SSC problem aims to partition $N$ data samples into $p$ groups such that similar data points are clustered together. In spectral clustering, a symmetric affinity matrix $W=[W_{ij}]_{N\times N}$ is constructed, where $W_{ij}\geq 0$ quantifies the pairwise similarity between samples $a_{i}$ and $a_{j}$. To promote sparsity and interpretability, the sparse spectral clustering framework is introduced in 2016ConvexSparseSpectralClustering , as follows
$\min_{X\in\mathrm{St}(N,p)}\langle L,XX^{\top}\rangle+\mu\|XX^{\top}\|_{1},$ (5)
where $\mu>0$ is the regularization parameter and $L=I_{N}-S^{-1/2}WS^{-1/2}$ is the normalized Laplacian matrix with $S^{1/2}$ being the diagonal matrix with diagonal elements $\sqrt{s_{1}},\sqrt{s_{2}},\dots,\sqrt{s_{n}}$ and $s_{i}=\sum_{j}W_{ij}$. It is not hard to see that problem (5) can be equivalently rewritten as:
$\min_{X\in\mathrm{St}(N,p),~{}Z\in\mathbb{R}^{N\times N}}\langle L,XX^{\top}\rangle+\mu\|Z\|_{1}+\mu\|XX^{\top}-Z\|_{1}.$ (6)
By the dual representation of the $\ell_{1}$-norm, i.e., $\|\cdot\|_{1}=\max_{\|Y\|_{\infty}\leq 1}\langle Y,\cdot\rangle$, problem (6) can be further equivalently reformulated as:
$\min_{X\in\mathrm{St}(N,p),~{}Z\in\mathbb{R}^{N\times N}}\max_{\|Y\|_{\infty}\leq\mu}\langle L,XX^{\top}\rangle+\mu\|Z\|_{1}+\langle Y,XX^{\top}-Z\rangle.$ (7)
Again, problem (7) is an instance of problem (1) with $\mathcal{M}$ being the product manifold $\mathrm{St}(N,p)\times\mathbb{R}^{N\times N}$, $S=\{Y\in\mathbb{R}^{N\times N}\mid\|Y\|_{\infty}\leq\mu\}$, $h(X,Z)=\mu\|Z\|_{1}$ and $f(X,Z,Y)=\langle L,XX^{\top}\rangle+\langle Y,XX^{\top}-Z\rangle$.
In recent years, nonconvex-concave (NC-C) minimax problems have received a tremendous amount of attention as they have found broad applications in diverse areas. Many studies focus on NC-C problems in the setting where both the constraints on $x$ and $y$ are closed convex sets and numerous methods have been developed, see, for example, linGradientDescentAscent ; linOptimalAlgorithmsMinimax ; nouiehedSolvingClassNonConvex ; ostrovskiiEfficientSearchFirstOrder2021 ; panEfficientAlgorithmNonconvexlinear2021a ; 2022Zeroth ; thekumparampilEfficientAlgorithmsSmooth ; xuUnifiedSingleloopAlternating2023 ; yangFasterSingleloopAlgorithms ; zhangSingleLoopSmoothedGradient . In contrast, there are relatively few works studying NC-C minimax problems in the setting where the constraint on $x$ is a Riemannian submanifold and the constraint on $y$ is a closed convex set. Among these works, Huang and Gao huangGradientDescentAscent2023 propose a Riemannian gradient descent-ascent algorithm for a smooth nonconvex-strongly concave (NC-SC) minimax problems and establish its iteration complexity of $\mathcal{O}(\varepsilon^{-2})$ for achieving $\varepsilon$-stationarity. Later, Xu et al. xuEfficientAlternatingRiemannian2024 present an alternating Riemannian/projected gradient descent-ascent (ARPGDA) algorithm with an iteration complexity of $\mathcal{O}(\varepsilon^{-3})$ for finding an $\varepsilon$-stationary point of smooth Riemannian nonconvex-linear (NC-L) minimax problem. Very recently, for problem (1) with additional assumptions that $h$ is smooth (but not necessarily convex) and $f(x,y)$ is linear with respect to $y$ for fixed $x$, the work xuRiemannianAlternatingDescent2024 extend the technique in and develop algorithms that can reach $\varepsilon$-stationary points within $\mathcal{O}(\varepsilon^{-3})$ iterations. It is worth noting that none of the methods in the aforementioned literature huangGradientDescentAscent2023 ; xuRiemannianAlternatingDescent2024 ; xuEfficientAlternatingRiemannian2024 can directly tackle problem (1) due to the following two reasons: (i) the convex function $h$ may be nonsmooth, and (ii) the function $f(x,y)$ is neither linear nor strongly concave with respect to $y$ for fixed $x$. Based on the needs of applications and the above background, it is highly desirable to develop efficient methods with theoretical convergence guarantees for solving the Riemannian nonsmooth NC-C minimax problem (1).
The main contributions of this paper are summarized below. First, we introduce the notions of optimization stationarity point as well as game stationarity for problem (1) and carefully study their relationships. Second, we propose a manifold proximal gradient descent ascent method for solving problem (1). At each iteration, this method performs one or multiple manifold proximal gradient descent steps to update $x$ and one proximal ascent step to update $y$. Third, we prove that the proposed method can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point within $O(\varepsilon^{-3})$ iterations. This iteration complexity result matches the state-of-the-art complexity analysis for single-loop algorithms in solving general nonsmooth NC-C optimization problems, where the constraint on $x$ is a compact convex set rather than a Riemannian manifold. Finally, we present numerical results on FSPCA and SSC to demonstrate the effectiveness of our proposed method.
The rest of the paper is organized as follows. Section 2 presents preliminaries that will be used in this paper. Section 3 discusses the optimality condition of problem (1). Section 4 introduces the proposed MPGDA method and establish its iteration complexity. Section 5 reports the numerical experiments.
2 Preliminaries
[…]
2.1 Riemannian submanifold
We first review the definition of an embedded submanifold in a Euclidean space $\mathbb{E}$. We say that function $\psi:U\subset\mathbb{E}\to\mathbb{R}^{k}$ is smooth if it is infinitely differentiable on $U$. Let $\mathrm{D}\psi(x)$ denote its differential at $x$.
Definition 1.
(boumal2023intromanifolds, , Definition 3.10) Let $\mathbb{E}$ be a Euclidean space of dimension $d$. A nonempty subset $\mathcal{M}$ of $\mathbb{E}$ is called an embedded submanifold of $\mathbb{E}$ of dimension $n$ if either $n=d$ and $\mathcal{M}$ is open in $\mathbb{E}$ or $n=d-k$ for some $k\geq 1$ and, for each $x\in\mathcal{M}$, there exists a neighborhood $U$ of $x$ in $\mathbb{E}$ and a smooth function $\psi:U\to\mathbb{R}^{k}$ such that:
  1. 1.
    If $y$ is in $U$, then $\psi(y)=0$ if and only if $y\in\mathcal{M}$; and
  2. 2.
    $\mathrm{rank}(\mathrm{D}\psi(x))=k$.
Such a function $\psi$ is called a local defining function for $\mathcal{M}$ at $x$.
Let $\mathcal{M}$ be an embedded submanifold of $\mathbb{E}$. The tangent space $T_{x}\mathcal{M}$ to $\mathcal{M}$ at $x\in\mathcal{M}$ is identified with $\mathbb{E}$ when $\mathcal{M}$ is open in $\mathbb{E}$; otherwise, $T_{x}\mathcal{M}=\mathrm{ker}(\mathrm{D}\psi(x))$ , where $\psi$ is a local defining function for $\mathcal{M}$ at $x$ (boumal2023intromanifolds, , Theorem 3.15). The embedded submanifold $\mathcal{M}$ is called a Riemannian submanifold of $\mathbb{E}$, when equipped with the Riemannian metric induced by the ambient space $\mathbb{E}$. Namely, for each $x\in\mathcal{M}$, the inner product on $T_{x}\mathcal{M}$ is defined as $\langle u,v\rangle_{x}:=\langle u,v\rangle$ for $u,~{}v\in T_{x}\mathcal{M}$. The normal space $N_{x}\mathcal{M}$ of a Riemannian submanifold $\mathcal{M}$ of $\mathbb{E}$ is the orthogonal complement of the tangent space $T_{x}\mathcal{M}$ in $\mathbb{E}$ (boumal2023intromanifolds, , Definition 5.48).
The tangent bundle of $\mathcal{M}$ is the disjoint union of its tangent spaces : $T\mathcal{M}:=\{(x,v)\mid x\in\mathcal{M},~{}v\in T_{x}\mathcal{M}\}$. A retraction on a manifold $\mathcal{M}$ is a smooth mapping $R$ from the tangent bundle $T\mathcal{M}$ onto $\mathcal{M}$ such that, for any $(x,v)\in T_{x}\mathcal{M}$, its restriction $R_{x}:T_{x}\mathcal{M}\to\mathcal{M}$ satisfies (i) $R_{x}(0)=x$ and (ii) $\mathrm{D}R_{x}(0)[v]=v$ (boumal2023intromanifolds, , Definition 3.47). The concept of a retraction offers a practical link between manifold and its tangent bundle. The following properties of retractions are needed for convergence analysis in this paper.
Proposition 1.
[…]
$\left\|R_{x}(v)-x\right\|$ $\leq M_{1}\left\|v\right\|,$
$\left\|R_{x}(v)-(x+v)\right\|$ $\leq M_{2}\left\|v\right\|^{2}.$
2.2 Generalized subdifferential
For a fucntion $\varphi:\mathbb{E}\to\mathbb{R}$, the usual one-sided directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined by
$\varphi^{\prime}(x;v):=\lim_{t\downarrow 0}\frac{\varphi(x+tv)-\varphi(x)}{t}.$
In contrast to the usual one-sided directional derivative, the (Clarke) generalized directional derivative dose not presuppose the existence of any limit (1983Optimization, , Section 2.1). Specifically, the (Clarke) generalized directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined as
$\varphi^{\circ}(x;v):=\lim\sup_{\begin{smallmatrix}y\to x\\ t\downarrow 0\end{smallmatrix}}\frac{\varphi(y+tv)-\varphi(y)}{t},$
and the (Clarke) generalized subdifferential of $\varphi$ at $x$ is defined as
$\partial\varphi(x):=\{z\in\mathbb{E}\mid\langle z,v\rangle\leq\varphi^{\circ}(x;v),~{}\forall v\in\mathbb{E}\}.$
The function $\varphi$ is said to be regular at $x$, provided that for all $v\in\mathbb{E}$, the usual one-sided directional derivative $\varphi^{\prime}(x;v)$ exists and $\varphi^{\prime}(x;v)=\varphi^{\circ}(x;v)$. If $\varphi_{1}$ and $\varphi_{2}$ are regular at $x$, then $\partial(\varphi_{1}+\varphi_{2})(x)=\partial\varphi_{1}(x)+\partial\varphi_{2}(x)$ (1983Optimization, , Definition 2.3.4). If $\varphi$ is a continuously differentiable function or a real-valued convex function, then it is regualr at any $x\in\mathbb{E}$ (1983Optimization, , Proposition 2.3.6). In addition, the indication function $\partial\delta_{\mathcal{M}}(x)$ of a Remannian submanifold $\mathcal{M}$ in $\mathbb{E}$ is regular at any $x\in\mathcal{M}$ (1983Optimization, , Theorem 2.4.7) and $\partial\delta_{\mathcal{M}}(x)=N_{x}\mathcal{M}$ (1983Optimization, , Proposition 2.4.12).
3 Optimality Conditions
In this section, we introduce two important concepts of stationarity for problem (1), namely, the optimization stationarity and game stationarity, and examine their relationships. To this end, we first definite the value function $\Phi:\mathbb{E}\to\mathbb{R}$ by
$\Phi(x):=\max_{y\in S}f(x,y)-g(y).$ (8)
Then the minimax problem (1) can be equivalently reformulated into
$\min_{x\in\mathcal{M}}\Phi(x)+h(x).$ (9)
The next proposition concerns the value function $\Phi$. Recalled that, a function $\varphi:\mathbb{E}\to\mathbb{R}$ is said to be $\ell$-weakly convex if $\varphi(\cdot)+(\ell/2)\|\cdot\|^{2}$ is convex.
Proposition 2.
Let $y_{x}\in\arg\max_{y\in S}\{f(x,y)-g(y)\}$. The value function $\Phi$ is $L_{x}$-weakly convex and for all $x\in\mathcal{M}$, it holds that $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$. Moreover, if $f(x,y)-g(y)$ is strongly concave in $y$, then $\Phi$ is differentiable and $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$.
Proof.
Because $f(x,y)$ is $L_{x}$-smooth respect to $x$, $f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex in $x$ for any given $y\in S$. Thus, $\Phi(x)+\frac{L_{x}}{2}\left\|x\right\|^{2}=\max_{y\in S}f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex. Therefore, $\Phi(x)$ is $L_{x}$-weakly convex. Since $S$ is compact, according to Danskin’s theorem (bernhardTheoremDanskinApplication1995, , Theorem D1), $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$, and in particular, $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$ when $f(x,y)-g(y)$ is strongly concave in $y$. ∎
In view of Proposition 2 and (1983Optimization, , Proposition 2.4.3), we see that any local minimizer $x^{*}$ of problem (9) must satisfy
$0\in\partial\Phi(x^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}).$ (10)
This naturally leads to the following definition of optimization stationarity.
Definition 2 (Optimization stationarity).
Consider problem (1) and the value function $\Phi$ defined in (8).
  1. (i)
    We say $x^{*}$ is a optimization-stationary point of problem (1) if it satisfies (10).
  2. (ii)
    Let $\varepsilon>0$ be given. Inspired by perez2019subdifferential , we say that $x^{*}$ is an $\varepsilon$-optimization-stationary point of problem (1) if
    $\mathrm{dist}(0,\partial_{\varepsilon}(\Phi+h)(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}))<\varepsilon,$
    with
    $\partial_{\varepsilon}(\Phi+h)(x^{*}):=\{z\in\mathbb{E}\mid\langle z,v\rangle-\varepsilon\leq(\Phi+h)^{\circ}(x^{*};v),~{}\forall v\in\mathbb{E}\}.$
Now we consider defining the game stationarity of problem (1). It is routine to prove that any (local) saddle point $(x^{*},y^{*})$ of problem (1) satisfies
$0\in\nabla_{x}f(x^{*},y^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}),$ (11a)
$0\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*}),$ (11b)
which yields the following definition of a game-stationary point of problem (1).
Definition 3 (Game-stationary point).
We say that $(x^{*},y^{*})$ is a game-stationary point of problem (1) if it satisfies (11a) and (11b).
Next, we study the relationships between optimization-stationary points and game-stationary points of problem (1).
Proposition 3.
If $x^{*}$ is an optimization-stationary point of the problem (1), then there exists $y^{*}\in S$ such that $(x^{*},y^{*})$ is a game-stationary point. Conversely, if $(x^{*},y^{*})$ is a game-stationary point of problem (1) , then $x^{*}$ is an optimization-stationary point.
Proof.
Suppose that $x^{*}$ is an optimization-stationary point of problem (1). By Proposition (2), $\Phi(x)+h(x)+L_{x}\|x-x^{*}\|^{2}$ is strongly convex. As $0\in\partial\delta_{\mathcal{M}}(x^{*})+\partial\Phi(x^{*})+\partial h(x^{*})$ and $\partial\delta_{\mathcal{M}}(x^{*})=N_{x^{*}}{\mathcal{M}}=\partial\delta_{T_{x^{*}}{\mathcal{M}}}(0)$, then $x^{*}$ is the unique minimum point of the problem
$\min_{x\in A_{x^{*}}\mathcal{M}}\Phi(x)+h(x)+L_{x}\|x-x^{*}\|^{2}$
with $A_{x^{*}}\mathcal{M}:=\{x\mid x=x^{*}+v,~{}v\in T_{x^{*}}{\mathcal{M}}\}$. Therefore, by Sion minimax theorem sionGeneralMinimaxTheorems1958 , there exists $y^{*}\in S$ such that
$f(x^{*},y^{*})+h(x^{*})-g(y^{*})$ $=\inf_{\begin{subarray}{c}x\in A_{x^{*}}{\mathcal{M}},\\ \|x-x^{*}\|\leq 1\end{subarray}}\sup_{y\in S}f(x,y)+h(x)-g(y)+L_{x}\|x-x^{*}\|^{2}$
$=\sup_{y\in S}\inf_{\begin{subarray}{c}x\in A_{x^{*}}{\mathcal{M}},\\ \|x-x^{*}\|\leq 1\end{subarray}}f(x,y)+h(x)-g(y)+L_{x}\|x-x^{*}\|^{2},$
which implies
$0\in\nabla_{x}f(x^{*},y^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}),$
$0\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*}).$
Conversely, suppose that $(x^{*},y^{*})$ is a game stationary point of problem (1). Then, according to Proposition 2, $y^{*}=\arg\max_{y\in S}f(x^{*},y)-g(y)$ and $\nabla_{x}f(x^{*},y^{*})\in\partial\Phi(x^{*})$. Thus, $0\in\partial\Phi(x^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*})$. ∎
Next, we introduce the game stationarity measure. For $(x,y)\in\mathbb{E}_{1}\times\mathbb{E}_{2}$, the game-stationarity measure is defined as
$\mathcal{G}^{\beta}(x,y):=\max\{\|\beta u\|,~{}\mathrm{dist}(0,\nabla_{y}f(x,y)-\partial g(y)-\partial\delta_{S}(y))\},$ (12)
where $\beta>0$ and $u=\arg\min_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$. This definition is motivated in part by the stationarity measure used in Euclidean minimax problem luHybridBlockSuccessive2020 . It is straightforward to verify that $(x^{*},y^{*})$ is a game-stationary point of problem (1) if and only if there exists $\beta>0$ such that $\mathcal{G}^{\beta}(x^{*},y^{*})=0$. Using this measure, we can also introduce the notion of $\varepsilon$-game-stationary point.
Definition 4.
Let $\varepsilon>0$ be given. We say $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1) if there exists $\beta>0$ such that $\mathcal{G}^{\beta}(x^{*},y^{*})\leq\varepsilon$.
Finally, we examine the connections between $\varepsilon$-optimization-stationary points and $\varepsilon$-game-stationary points of problem (1).
Proposition 4.
Let $\varepsilon>0$ be given. Suppose that $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1) with $\mathcal{G}^{\beta}(x^{*},y^{*})\leq\varepsilon$. Then $x^{*}$ is a $(C\varepsilon)$-optimization-stationary poin, where $C=(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}/\beta$ and $\sigma_{y}:=\sup_{y\in S}\left\|y\right\|$.
Proof.
Suppose the point $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1), then there exists $z$ such that $\|z\|=1$ and $\varepsilon z\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*})$. For $u=\arg\min_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$, let
$\Phi_{\varepsilon}^{u}(x):=$ $\max_{y\in S}f(x+u,y)-g(y)-\frac{\varepsilon}{2}\|y-(y^{*}-z)\|^{2}$
$+\langle\nabla_{y}f(x^{*},y^{*})-\nabla_{y}f(x^{*}+u,y^{*}),y\rangle.$
By Proposition 2 and the choice of $z$, we have $\nabla\Phi_{\varepsilon}^{u}(x^{*})=\nabla_{x}f(x^{*}+u,y^{*})$ and
$\Phi_{\varepsilon}^{u}(x^{*})+\langle\nabla_{x}f(x^{*}+u,y),x-(x^{*}+u)\rangle\leq\Phi_{\varepsilon}^{u}(x-u)+\frac{L_{x}}{2}\|x-(x^{*}+u)\|^{2}.$
Because
$\Phi(x^{*}+u)-\varepsilon(2\sigma_{y}^{2}+1)-L_{y}\sigma_{y}\|u\|\leq\Phi_{\varepsilon}^{u}(x^{*})$
and
$\Phi_{\varepsilon}^{u}(x-u)\leq\Phi(x)+L_{y}\sigma_{y}\|u\|,$
we arrive at
$\Phi(x^{*}+u)+\langle\nabla_{x}f(x^{*}+u,y),x-(x^{*}+u)\rangle-\varepsilon(2\sigma_{y}^{2}+1)-2L_{y}\sigma_{y}\|u\|$
$\leq$ $\Phi(x)+\frac{L_{x}}{2}\|x-(x^{*}+u)\|^{2}.$
Since $h$ is convex and $L_{h}$-Lipschitz continuous, then
$h(x^{*}+u)+\langle\xi,x-(x^{*}+u)\rangle\leq h(x)$
for all $\xi\in\partial h(x^{*}+u)$. Therefore, by combing the above two inequalities, we obtain
$\nabla_{x}f(x^{*}+u,y^{*})+\partial h(x^{*}+u)\subset\partial_{\hat{\varepsilon}}(\Phi+h)(x^{*}+u),$
with $\hat{\varepsilon}=\varepsilon(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}\|u\|$. Furthermore, as $\|\mathcal{G}_{x}^{\beta}(x^{*},y^{*})\|\leq\varepsilon$, we can conclude that $x^{*}$ is a $(C\varepsilon)$-optimization-stationary point with $C=(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}/\beta$. ∎
4 The proposed MPGDA algorithm
In this section, we propose our MPGDA method for solving problem (1) and establish its iteration complexity. Our algorithm development is partially motivated by heApproximationProximalGradient2024 ; luHybridBlockSuccessive2020 ; xuRiemannianAlternatingDescent2024 . For the $k$-th iteration, we introduce the following value function:
$Q_{k}(x):=h(x)+\Phi_{k}(x)$ (13)
where $\Phi_{k}:\mathbb{E}_{1}\to\mathbb{R}$ is defined at $x\in\mathbb{E}_{1}$ by
$\Phi_{k}(x):=\max_{y\in S}\{f(x,y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}\}.$ (14)
Here, $y_{k}$ denotes the iterate with respect to $y$ obtained in the previous iteration, $\gamma_{k}>0$ is the regularization parameter, and $\rho_{k}>0$ is the proximal parameter. Moreover, we choose $\gamma_{k}=\gamma_{0}/k^{1/3}$ with some $\gamma_{0}>0$ and assume the non-increasing sequence $\{\rho_{k}\}$ to be summable, i.e.,
$\{\rho_{k}\}\in\mathcal{S}:=\{\{\rho_{k}\}\mid\sum_{k=0}^{\infty}\rho_{k}<+\infty,~{}~{}\rho_{k}\geq\rho_{k+1}>0,k=0,1,\dots~{}\}.$
We further introduced the mapping $\bar{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ defined by
$\bar{y}_{k}(x):=\arg\max_{y\in S}\{f(x,y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}\}.$ (15)
Then it holds that $Q_{k}(x)=F(x,\bar{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x)-y_{k}\|^{2}$ and in view of Danskin’s theorem (danskinTheoryMaxMinApplications1966, , Theorem D1), $\Phi_{k}$ is smooth on $\mathbb{E}_{1}$ and $\nabla\Phi_{k}(x)=\nabla_{x}f(x,\bar{y}_{k}(x))$.
Now, at the $k$-th iteration, the proposed MPGDA method computes the new iterate by approximately solving the subproblem:
$\min_{x\in\mathcal{M}}Q_{k}(x).$ (16)
Inspired by the structure of $Q_{k}$ and the ManPG algorithm (chenProximalGradientMethod2020, , Algorithm 4.1), we perform Riemannian manifold proximal gradient descent steps for $T_{k}$ times to generate $x_{k+1}$ with $T_{k}$ being an integer in $[1,\overline{T}]$ where $\overline{T}$ is a predetermined positive integer. Then $y_{k+1}$ is updated as $\bar{y}_{k}(x_{k+1})$ with $\bar{y}_{k}$ defined in (15). Notice that the update of $y$ can also be interpreted as performing a proximal ascent step on the function $f(x_{k+1},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}$. We formally presented the proposed method in Algorithm 1.
Algorithm 1 Manifold Proximal Gradient Descent Ascent (MPGDA) for problem (1)
0:  $c_{1}>0$, $\eta\in(0,1)$, $\gamma_{0}>0$, $x_{0}\in\mathcal{M}$, $y_{0}\in\mathrm{dom}(g)\cap S$, $0<l_{\min}<l_{\max}$, $\{\rho_{k}\}\in\mathcal{S}$, $\{\gamma_{k}=\gamma_{0}/k^{1/3}:k\in N\}$, $\{T_{k}\}\subset[1,\overline{T}]$;
1:  for $k=0,\dots$ do
2:     Set $x_{k,0}=x_{k}$,
3:     for $i=0,\dots,T_{k}-1$ do
4:        Choose $l_{k,i}\in[l_{\min},l_{\max}]$ and set $\beta_{k,i}=l_{k,i}/(\rho_{k}+\gamma_{k})$.
5:        Compute
$v_{k,i}:=\arg\min_{v\in T_{x_{k,i}}\mathcal{M}}\langle\nabla_{x}f(x_{k,i},\bar{y}_{k}(x_{k,i})),v\rangle+h(x_{k,i}+v)+\frac{\beta_{k,i}}{2}\left\|v\right\|^{2},$ (17)
6:        Find the smallest non-negative integer $j$ such that
$Q_{k}(R_{x_{k,i}}(\eta^{j}v_{k,i}))\leq Q_{k}(x_{k,i})-c_{1}\eta^{j}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$ (18)
and update $x_{k,i+1}=R_{x_{k,i}}(\eta^{j}v_{k,i})$
7:     end for
8:     $x_{k+1}=x_{k,T_{k}},~{}y_{k+1}=\bar{y}_{k}(x_{k+1})$.
9:  end for
10:  return $(x_{k+1},y_{k+1})$.
Before proceeding, we make two remarks on the iteration procedure of MPGDA algorithm. First, given $x_{k,i}\in\mathcal{M}$, MPGDA algorithm computes a descent direction $v_{k,i}$ of $Q_{k}$ restricted to the tangent space $T_{x_{k,i}}\mathcal{M}$ via tackling the linear constrained strongly convex optimization problem (17), which can usually be efficiently solved, see (li2024proximal, , Section 4), for example. Second, in practice, we make some choices of parameters $l_{k,i}$ and $\rho_{k}$, which enhance the efficiency of MPGDA algorithm in our test and also satisfy $l_{k,i}\in[l_{\min},l_{\max}]$ and $\rho_{k}\in S$. For parameters $l_{k,i}$, we adopt the Riemannian Barzilai-Borwein (BB) stepsize iannazzo2018riemannian ; 2013AWen as follows
$l_{k,i}=\begin{cases}\min\{\max\{l_{\min},(\rho_{k}+\gamma_{k})|\frac{\langle\Delta X,\Delta G\rangle}{\|\Delta X\|^{2}}|)\},l_{\max}\},&if~{}\Delta X\neq 0,\\ l_{\max},&if~{}\Delta X=0,\end{cases}$
where $\Delta X=X_{k,i}-X_{k,i-1}$ and $\Delta G=\mathrm{grad}\Phi_{k}(x_{k,i})-\mathrm{grad}\Phi_{k}(x_{k,i-1})$. Moreover, inspired by xuRiemannianAlternatingDescent2024 , for given constant $\theta>1$, $\tau_{1}\in(0,1)$ and $\tau_{2}\in(0,1)$, we set $\rho_{0}=\xi_{0}$ and for $k\geq 1$,
$\rho_{k}=\frac{\xi_{k}}{k^{\theta}}~{}~{}\text{with}~{}~{}\xi_{k}=\begin{cases}\tau_{2}\xi_{k-1},&\text{if}~{}~{}\delta_{k}\geq\tau_{1}\delta_{k-1},\\ \xi_{k-1},&\text{else},\end{cases}$ (19)
where $\delta_{k}$ is defined as
$\delta_{k}=\|\gamma_{k-1}y_{k}+\rho_{k-1}(y_{k}-y_{k-1})\|_{\infty}.$
In what follows, we conduct convergence analysis for MPGDA algorithm. First, we present a technical lemma that guarantees the boundedness of the mapping $\bar{y}_{k}$.
Lemma 1.
Let the mapping $\bar{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ be defined in (15) for $k\in N$. Then for all $x$, $\bar{x}\in\mathbb{E}_{1}$ and $y\in\mathbb{E}_{2}$ , it holds that
$f(x,y)-g(y)$
$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
$+\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x-\bar{x}\|^{2}.$
Proof.
According to the definition of $\bar{y}_{k}(\bar{x})$, we have
$0\in\nabla_{y}f(\bar{x},\bar{y}_{k}(\bar{x}))-\partial g(\bar{y}_{k}(\bar{x}))-\partial\delta_{S}(\bar{y}_{k}(\bar{x}))-\rho_{k}(\bar{y}_{k}(\bar{x})-y_{k})-\gamma_{k}\bar{y}_{k}(\bar{x}).$
By subgradient inequality, we obtain
$f(x,y)-g(y)$
$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))+\langle\nabla_{y}f(x,\bar{y}_{k}(\bar{x}))-\nabla_{y}f(\bar{x},\bar{y}_{k}(\bar{x})),y-\bar{y}_{k}(\bar{x})\rangle$
$+\langle\rho_{k}(\bar{y}_{k}(\bar{x})-y_{k})+\gamma_{k}\bar{y}_{k}(\bar{x}),y-\bar{y}_{k}(\bar{x})\rangle.$
Since $2\langle a,a-b\rangle=\left\|a\right\|^{2}+\left\|a-b\right\|^{2}-\left\|b\right\|^{2}$,by Cauchy-Schwarz inequality, we arrive at
$f(x,y)-g(y)$
$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
$+\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x-\bar{x}\|^{2}.$
Next, we prove that the descent condition (18) must be satisfied after a unified finite number of iterations for all $k\in N$ and $i\in[0,T_{k}-1]$, which ensures that the proposed Algorithm 1 is well-defined.
Proposition 5.
Consider Algorithm 1. There exists a constant $\bar{\alpha}>0$ such that for all $k\in N$ and $0<\alpha\leq\min\{1,~{}\bar{\alpha}\}$, it holds that
$Q_{k}(R_{x_{k,i}}(\alpha v_{k,i}))\leq Q_{k}(x_{k,i})-\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$
for all $i=0,1,\dots,T_{k}-1$.
Proof.
By (chenProximalGradientMethod2020, , Lemma 5.2), for any $0<\alpha\leq\min\{1,~{}\frac{\beta_{k,i}}{2M_{2}(L_{h}+G)+M_{1}^{2}L_{x}}\}$,
$f(x_{\alpha},\bar{y}_{k}(x_{k,i}))+h(x_{\alpha})\leq f(x_{k,i},\bar{y}_{k}(x_{k,i}))+h(x_{k,i})-\frac{\alpha\beta_{k,i}}{2}\|v_{k,i}\|^{2},$
where $x_{\alpha}=R_{x_{k,i}}(\alpha v_{k,i})$ and $G=\sup_{x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S}\|\nabla_{x}F(x,y)\|$.
According to Lemma 1, we have
$f(x_{\alpha},\bar{y}_{k}(x_{\alpha}))-g(\bar{y}_{k}(x_{\alpha}))$
$\leq f(x_{\alpha},\bar{y}_{k}(x_{k,i}))-g(\bar{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{k,i})\|^{2}+\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{\alpha})\|^{2}$
$+\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{\alpha})-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x_{\alpha}-x_{k,i}\|^{2}.$
By summing these inequalities with the inequality
$\|x_{\alpha}-x_{k,i}\|\leq\alpha M_{1}\|v_{k,i}\|$
in Proposition 1, we derive the following inequality:
$F(x_{\alpha},\bar{y}_{k}(x_{\alpha}))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{\alpha})\|^{2}$
$\leq F(x_{k,i},\bar{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{k,i})\|^{2}$
$+\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{\alpha})-y_{k}\|^{2}-(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}.$
Since $\beta_{k,i}=l_{k,i}/(\rho_{k}+\gamma_{k})$, for any $0<\alpha\leq\frac{l_{k,i}}{2M_{1}^{2}L_{y}^{2}}$, we have
$(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}$ $=\beta_{k,i}(\frac{\alpha}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2l_{k,i}})\|v_{k,i}\|^{2}\geq\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}.$
Note that $Q_{k}(x)=F(x,\bar{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x)-y_{k}\|^{2}$, then for any $0<\alpha\leq\min\{1,\frac{l_{\min}}{(\rho_{0}+\gamma_{0})(2M_{2}(L_{h}+G)+M_{1}^{2}L_{x})},\frac{l_{\min}}{2M_{1}^{2}L_{y}^{2}}\}$, we have
$Q_{k}(x_{\alpha})\leq Q_{k}(x_{k,i})-\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$
for all $k\geq 0$. ∎
We are now ready to establish the iteration complexity for MPGDA algorithm. Let $\varepsilon>0$ be a given target accuracy, we define
$T(\varepsilon):=\min\{k\mid\mathcal{G}^{\beta_{k,0}}(x_{k,0},\bar{y}_{k}(x_{k,0}))<\varepsilon\}$
where $\mathcal{G}^{\beta}(\cdot,\cdot)$ is defined in (12). In other words, $T(\varepsilon)$ stands for the minimal number of outer iterations needed for MPGDA algorithm to obtain a $\varepsilon$-game-stationary point. We further define two constants $\overline{Q}$ and $\underline{Q}$ as
$\overline{Q}$ $:=\sup\{{F(x,y)\mid x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S}\},$
$\underline{Q}$ $:=\inf\{F(x,y)-\frac{\rho_{0}}{2}\|y-y_{0}\|^{2}-\frac{\gamma_{0}}{2}\|y\|^{2}\mid x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S\}.$
Theorem 4.1.
Given $\varepsilon>0$, let $\{(x_{k},y_{k})\}$ be generated by Algorithm 1. Then, there exists a positive $\bar{\alpha}>0$ such that the following inequality holds:
$T(\varepsilon)\leq\frac{\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\}}{\varepsilon^{3}}$
with
$C=\frac{\sqrt{2(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2})l_{\max}}}{\sqrt{3\bar{\alpha}\eta\gamma_{0}}}.$
Here, all the parameters $\rho_{0},~{}\gamma_{0},~{}\sigma_{y},~{}\overline{Q},~{}\underline{Q},~{}\overline{T},~{}S_{\rho},~{}l_{\max},~{}\eta$ are defined as in the relevant context of the paper.
Proof.
Let $\beta_{k}:=\beta_{k,0},~{}v_{k}:=v_{k,0}$. Since $\gamma_{k+1}\leq\gamma_{k}$ and $\rho_{k+1}\leq\rho_{k}$, by the definition of $Q_{k}$, we have
$Q_{k+1}(x_{k+1})$ $=h(x_{k+1})+\max_{y\in S}f(x_{k+1},y)-g(y)-\frac{\gamma_{k+1}}{2}\|y\|^{2}-\frac{\rho_{k+1}}{2}\|y-y_{k+1}\|^{2}$
$\leq h(x_{k,T_{k}})+\max_{y\in S}f(x_{k,T_{k}},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+4\rho_{k}\sigma_{y}^{2}$
$=Q_{k}(x_{k,T_{k}})+4\rho_{k}\sigma_{y}^{2}.$
Furthermore, according to Proposition 5, there exists $\bar{\alpha}>0$ such that
$Q_{k+1}(x_{k+1})\leq Q_{k}(x_{k})+2(T_{k}+2)\rho_{k}\sigma_{y}^{2}-\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}.$
Summing the above inequality over $k=K_{1},K_{1}+1,\dots,K$ yields
$\sum_{k=K_{1}}^{K}\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}\leq{Q}_{K_{1}}(x_{K_{1}})-{Q}_{K}(x_{K})+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2}$ (20)
with $S_{\rho}=\sum_{k=0}^{\infty}\rho_{k}<\infty$.
Since $\beta_{k}\leq\frac{l_{\max}}{\gamma_{0}k^{-1/3}}$, then
$\sum_{k=K_{1}}^{K}\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}\geq(\frac{\bar{\alpha}\eta\gamma_{0}}{l_{\max}}\sum_{k=K_{1}}^{K}k^{1/3})\min_{K_{1}\leq k\leq K}\|\beta_{k}v_{k}\|^{2}.$ (21)
Note that
[…]
let $K=\lceil{2^{3/2}K_{1}}\rceil$ , then we obtain
$\sum_{K_{1}\leq k\leq\lceil{2^{3/2}K_{1}}\rceil}\frac{1}{k^{1/3}}\geq\frac{3}{2}K_{1}^{2/3}.$ (22)
Therefore, combining (20), (21) and (22), with $Q_{K}(x_{K_{1}})\leq\overline{Q}$ and $Q_{\lceil{2^{3/2}K_{1}}\rceil}(x_{\lceil{2^{3/2}K_{1}}\rceil})\geq\underline{Q}$, we have
$\min_{K_{1}\leq k\leq\lceil{2^{3/2}K_{1}}\rceil}\|\beta_{k}v_{k}\|^{2}\leq\frac{2l_{\max}}{3\bar{\alpha}\eta\gamma_{0}K_{1}^{2/3}}(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2}).$ (23)
In view of definition (15) and the associated first-order optimality condition, for $\bar{y}_{k}(x_{k,0})$, we have that
$\rho_{k}(\bar{y}_{k}(x_{k,0})-y_{k})+\gamma_{k}\bar{y}_{k}(x_{k,0})\in\nabla_{y}f(x_{k,0},\bar{y}_{k}(x_{k,0}))-\partial g(\bar{y}_{k}(x_{k,0}))-\partial\delta_{S}(\bar{y}_{k}(x_{k,0})).$
Invoking this inclusion, we further obtain that
$\mathrm{dist}(0,\nabla_{y}f(x_{k,0},\bar{y}_{k}(x_{k,0}))-\partial g(\bar{y}_{k}(x_{k,0}))-\partial\delta_{S}(\bar{y}_{k}(x_{k,0})))$ (24a)
$\leq$ $\|\rho_{k}(\bar{y}_{k}(x_{k,0})-y_{k})+\gamma_{k}\bar{y}_{k}(x_{k,0})\|$ (24b)
$\leq$ $\sigma_{y}(2\rho_{k}+\gamma_{k}).$ (24c)
Finally, by the definition of $\rho_{k}$, $\gamma_{k}$ and $\mathcal{G}^{\beta_{k}}$ , combining (23) and (24), we conclude that
$\min_{K_{1}\leq k\leq\lceil 2^{3/2}K_{1}\rceil}\mathcal{G}^{\beta_{k}}(x_{k,0},\bar{y}_{k}(x_{k,0}))\leq\frac{\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\}}{K_{1}^{1/3}}$
with
$C=\frac{\sqrt{2(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2})l_{\max}}}{\sqrt{3\bar{\alpha}\eta\gamma_{0}}},$
and the proof is finished. ∎
Theorem 4.1 demonstrates that the proposed MPGDA algorithm can find an $\varepsilon$-game-stationary point for problem (1) within $O(\varepsilon^{-3})$ iterations. In view of Proposition 4, this also indicates that the iteration complexity of the MPGDA algorithm for returning an $\varepsilon$-optimization-stationary point is $\mathcal{O}(\varepsilon^{-3})$.
5 Numerical experiments
In this section, we demonstrate the efficiency of the proposed MPGDA algorithm by applying it to FSPCA problem (4) and SSC problem (7). All the experiments are implemented in MATLAB R2019a and conducted on a standard PC with 3.40GHz Intel(R) Core(TM) i5-7500 CPU and 16GB of RAM.
We first specify some implementation details of the proposed algorithm. Throughout the tests, we set $c_{1}=10^{-4}$, $\eta=0.1$, $l_{\min}=10^{-16}$, $l_{\max}=10^{16}$, $\tau_{1}=0.999$ and $\tau_{2}=0.9$. The parameters $\gamma_{0}$, $\xi_{0}>0$, $\theta>1$ and $\{T_{k}\}$ vary across the two application problems. Since $h$ is the $\ell_{1}$-norm in both problems, the semismooth Newton method developed in (chenProximalGradientMethod2020, , Section 4.2) is used to solve subproblem (17). We utilize the QR factorization as the retraction on the Stiefel manifold 2008Optimization .
5.1 Experiments on FSPCA
In this subsection, we compare the proposed MPGDA algorithm with the constraint relaxed majorization-minimization (CRMM) algorithm (babuFairPrincipalComponent2023, , Algorithm 3) for solving FSPCA problem (4). The CRMM algorithm first relaxes the semi-orthogonality constraint $X^{T}X=I$ into a linear matrix inequality constraint and then applies some majorization-minimization iteration procedure, which requires solving a quadratic program at each iteration.
We perform tests on synthetic datasets and a real dataset where the samples belong to two groups. The synthetic datasets are generated following similar settings to (2016SparseGeneralizedEigenvalueProblem, , Section 5.1). In detail, for the $i$-th group, we randomly generate $m_{i}=200$ samples with $d=40$ features according to Gaussian distribution with mean $\mu^{i}\in\mathbb{R}^{d}$ and covariance $\Sigma\in\mathbb{R}^{d\times d}$. We set $\mu_{j}^{1}=0,\ \mu_{j}^{2}=1/3$, for $j\in\{2,4,\cdots,200\}$, and $\mu_{j}^{i}=0$ otherwise. Meanwhile, the covariance matrix $\Sigma$ is chosen as a block diagonal matrix with five $8\times 8$ blocks, where the $(j,j^{\prime})$-th entry of each block is $0.8^{|j-j^{\prime}|}$. For the real-world data, we use the default credit dataset yeh2009creditdata and preprocess it as done in samadiPriceFairPCA . We set the regularization parameter $\mu=0.1$ in our tests. Furthermore, the CRMM algorithm is terminated if the relative error between the solutions obtained over two successive iterations is less than $10^{-6}$. For MPGDA algorithm, we set $\gamma_{0}=10^{-6}$, $\xi_{0}=4\sqrt{r}*10^{4}$, $\theta=1.5$, $T_{k}\equiv 15$, and terminate the algorithm if it satisfies:
$\mathcal{G}^{\beta_{k,0}}(x_{k},y_{k})<\varepsilon$ (25)
with $\varepsilon=10^{-6}$. Both compared algorithms start from the same random initial points on $\mathrm{St}(d,r)$ and the maximum iteration unmber of them are set as 1000.
We present the average results over 50 runs on different synthetic datasets and results for the default credit dataset in Table 1. We report the objective values (Obj), and CPU time in seconds (Time) and the number of iterations (Iter). It is observed that MPGDA algorithm achieves similar objective values to the CRMM algorithm while requiring much less computational time. Specially, our proposed algorithm is on average more than 20 times faster than the CRMM algorithm.
Table 1: Results for FSPCA
CRMM MPGDA
Dataset r Obj. Time (s) Iter. Obj. Time (s) Iter.
Synthetic 2 $-9.795$ $0.269$ $201$9.808 0.007 167
3 $-14.380$ $0.344$ 155 -14.382 0.010 $192$
4 $-18.714$ $0.443$ 162 -18.745 0.012 $217$
5 $-22.774$ $0.209$ 83 -22.809 0.008 $197$
Credit 2 $-9.531$ $4.014$ 93 -9.742 0.133 $120$
3 $-10.557$ $15.692$ $271$10.802 0.104 91
4 $-11.637$ $8.977$ $142$11.923 0.010 5
5 $-12.572$ $20.058$ $277$12.724 0.012 4
5.2 Experiments on SSC
In this subsection, we compare the proposed MPGDA method with the RADA-RGD and RADA-PGD algorithms in xuRiemannianAlternatingDescent2024 for solving the SSC problem. Note that in xuRiemannianAlternatingDescent2024 , the original SSC problem (5) is reformulated into the following smooth minimax problem on the Grassmann manifold so that the RADA-PGD and RADA-RGD algorithms can be directly applied:
$\min_{Z\in\mathrm{Gr}(N,p)}\max_{\|Y\|_{\infty}\leq\mu}\{\langle L,Z\rangle+\langle Y,Z\rangle\}.$ (26)
Here, $\mathrm{Gr}(N,p)=\{XX^{\top}\in\mathbb{R}^{N\times N}\mid X\in\mathrm{St}(N,p)\}$ denotes the Grassmann manifold AGrassmannmanifoldhandbook2024A . The codes of these two algorithms for solving (26) are downloaded from: https://github.com/XuMeng00124/RADAopt.
In the tests, we generate synthetic datasets as in 2016ConvexSparseSpectralClustering , where the data points $\{a_{i}\}_{i=1}^{N}$ are independently drawn from the standard Gaussian distribution and $W_{ij}=|\langle a_{i},a_{j}\rangle|$. For the two RADA algorithms, we adopt the parameters as suggested in (xuRiemannianAlternatingDescent2024, , Section 6.3). Specifically, they are terminated once an $\varepsilon$-game-stationary point of problem (26) is found with $\varepsilon=10^{-4}$ and the initial point is set as $X_{0}X_{0}^{\top}$, where $X_{0}$ consists of $p$ eigenvectors associated with the $p$ smallest eigenvalues of $L$. For MPGDA algorithm, we set $\gamma_{0}=10^{-5}$, $\xi_{0}=\sqrt{p}N^{2}$, $\theta=2$ and $T_{k}\equiv 3$. Moreover, we choose the aforementioned $X_{0}$ as the initial point of MPGDA algorithm and terminate the method if it satisfies (25) with $\varepsilon=10^{-4}$. The maximum number of iterations for all the compared algorithms is set as 1000.
Table 2: Results for SSC
Config Para $(p/\mu)$ RADA-PGD RADA-RGD MPGDA
Obj. Time Iter. Obj. Time Iter. Obj. Time Iter.
$\mu=0.1$ $N=200$ 2 2.187 $0.513$ $265$ 2.187 $0.796$ $153$ 2.187 0.114 65
4 $4.374$ $0.530$ $259$ 4.373 $0.812$ $157$ $4.374$ 0.116 64
6 6.560 $0.521$ $224$ 6.560 $0.855$ $164$ 6.560 0.117 63
8 8.747 $0.580$ $198$ 8.747 $0.861$ $166$ 8.747 0.129 64
10 10.934 $0.621$ $180$ 10.934 $0.995$ $167$ 10.934 0.129 65
$p=5$ $N=200$ 0.1 5.467 $0.535$ $244$ 5.467 $0.868$ $164$ 5.467 0.122 63
0.2 5.967 $0.556$ $263$ 5.967 $0.934$ $178$ 5.967 0.133 71
0.5 7.467 $0.593$ $286$ 7.467 $1.071$ $202$ 7.467 0.174 86
1.0 9.967 $0.616$ $300$ 9.967 $1.188$ $228$ 9.967 0.205 95
The computational results averaged over 50 synthetic data sets with $N=200$ data points are summarized in Table 2, where we report the obtained objective function value of problem (5) (Obj.), the CPU time in seconds (Time), and number of outer iterations (Iter.). We observe that all the compared algorithms attain comparable objective function value for different number of groups and regularization parameters $\mu$. In addition, the proposed MPGDA method consistently outperform the two RADA algorithms in terms of CPU time, which again validates its computational efficiency.